# Building on Efficient Foundations: Effectively Training LLMs with Structured Feedforward Layers

### v1

@@ 4.2 Findings on efficient training
Figure 6: Scaling curves of structured matrices: The scaling curves during training between the standard Transformer and the modified version with a Structured FFN, which retains 63% and 32% of the original parameters, respectively.
Here, we analyze the performance of Transformers using FFN with 63% or 32% of the original FFN parameters by reducing the matrix rank and number of blocks. Experiments are conducted with various Transformer sizes, comparing performance among different structured matrices and showing the scaling curves.
Comparison between structured FFNs We conducted experiments fixing the model and training FLOPs for LowRank and BlockDense when comparing them to BlockShuffle in Sec. 4 and Table 9. We achieved a 0.8 perplexity reduction on Transformer-s and a 0.4 perplexity reduction on Transformer-m. We think that for FFNs in language models, BlockShuffle may not be an ideal option. However, in vision tasks where ShuffleNet has demonstrated success and block-diagonal matrices are better for local information, we show in Sec. B.3 that the block-diagonal matrix is a more suitable inductive bias. Moreover, LowRank and BlockDense, with 63% of the parameters in FFN, increase the loss by only 0.02 to 0.04 while reducing the total training time by about 15% on Transformer-xl and Transformer-l. Additionally, they can accelerate training by 1.36 times with only a 1 perplexity increase on Transformer-xl with 32% of the parameters.
Scaling analysis We compare the performance of different model sizes and show their scaling curves in 1(a), 6(a), 6(b). All the efficient parametrization gets closer to the baseline with the increasing model size. For example, LowRank with 32% parameters has a $0.08$ loss gap to Transformer-xl compared to a loss gap of $0.12$ at the scale of Transformer-s models.
According to 1(a), we observe that: (i) The structured matrices exhibit steeper scaling curves compared to the dense networks, indicating significant potential for these efficient designs in LLMs. (ii) The scaling curve of 32% parameters of FFN is steeper than the 63% parameters of FFN highlights the scaling potential of highly structured large models. (iii) Given fixed training FLOPs budget, a wider and structured network with more tokens may achieve comparable or superior performance to dense networks at the optimal trade-off.
Our scaling curves for the structured matrices can be further optimized: (1) they are not drawn at their optimal training-compute trade-off while the baseline is. (2) Only the FFN is made structured, while attention remains dense and thus contributes more to the whole model, which is less optimal. The second point also explains why the current 32% parameter curve shows a larger validation loss than the 63% parameter curve and motivates us to reduce attention further using existing techniques in Sec. 4.4.
@@ 4.3 Self-guided training
We apply self-guided training during the first half of training to demonstrate its effectiveness. As shown in Sec. 4.3 and Sec. B.2, our method consistently reduces loss across all efficient parametrizations, improving the perplexity by 1.2 for Transformer-s and 0.8 for Transformer-m. Then, to enable a straightforward comparison under the same training FLOPs, we adjust the training steps for self-guided training and repeat those tokens at the end to ensure they’re fully learned by structured matrices. As can be seen in Sec. B.5 and 12(a), 1(b), 12(b), this reduces the perplexity gap for Transformer-xl from 1.0, 1.2, and 1.3 to 0.5, 0.5, and 0.6 for LowRank, BlockDense, and BlockShuffle, respectively, under the same training FLOPs and 32% inference FFN FLOPs. Additionally, we compare our method with another advanced baseline that trains structured parametrizations with more tokens, showing that ours achieves comparable or superior results even with the same number of tokens.
Table 2: Wide and structured network performance: For GQA, we increase the intermediate FFN size as in [6] to match parameters. For ours, we either use more tokens or self-guided training (marked with ) to match the training FLOPs. TP refers to the maximum throughput measured on a generation length of 256. Detailed configuration is put in Table 12.
@@
Table 3: Ablation study about self-guided training across methods. We apply it to the first half of training with 25% increased FFN budget. For more results, please refer to Sec. B.2.
4.4 Wide and structured network
In the previous section, we explored structuring FFN, achieving steeper scaling curves and better results via self-guided training. As maintaining the parameter ratio of attention to FFN can be important, in this section, we use GQA [8] to make attention efficient and LowRank for FFN, designing a wide and structured network from Transformer-m and Transformer-l (see Table 2). To match the training FLOPs, we either train on more tokens or apply self-guided training. This approach aims to enhance efficiency with a much smaller network, achieving an 8% and 17% maximum throughput boost on Transformer-m and Transformer-l, respectively, while maintaining or slightly improving perplexity.
5 Conclusion
In this paper, we conducted extensive experiments investigating the use of structured matrices to parameterize FFN in Transformers, with models up to 1.3B parameters on the RefinedWeb dataset. Our primary aim was not to determine which structured matrices perform best, as this can be task-dependent, but to explore common issues including efficiency and optimization challenges of existing structured matrices as well as BlockDense. For example, with the proposed self-guided training, our models can achieve a 1.36× inference speed-up with only a 0.4 perplexity increase. With wide and structured networks, we show that we can improve both throughput and perplexity.
Limitations: BlockDense and BlockShuffle are more complicated than LowRank. In this work, we only explored a limited range of hyperparameter settings of them. However, since these approaches are new, we believe that further performance improvements may be possible by better tuning their hyperparameters. We primarily focused on language modeling with limited vision experiments included in the appendix. Additionally, we did not explore the optimal scaling laws for structured matrices, which may further enhance performance. We hope this work will encourage further exploration of other structured linear parameterizations.
Acknowledgments and Disclosure of Funding
We are grateful to Soham De for the insightful discussions and his valuable feedback on our work.
Appendix
Appendix A Supplementary of optimization challenges
A.1 Observation of training dynamics
We provide additional training dynamics curves, including BlockDense, BlockShuffle in Fig. 8 and Fig. 9, highlighting their loss spikes or slower convergence. In contrast, we show the improved loss curves and dynamics from applying self-guided training to the worst case.
@@ A.1 Observation of training dynamics
Meanwhile, we report the loss spikes found on the Transformer-xl and RefinedWeb dataset in Fig. 10.
@@
A.2 On the poorer training dynamics
$\Delta{\bm{W}}={\bm{W}}^{\prime}-{\bm{W}}=-lr\cdot{\bm{g}}{\bm{x}}^{\top}.$ (3)
And the one for ${\bm{U}}{\bm{V}}$:
Thus, it can be seen from Eq. (missing) 4 that the projections ${\bm{U}}{\bm{U}}^{\top}$ and ${\bm{V}}^{\top}{\bm{V}}$ can disrupt the gradient ${\bm{g}}{\bm{x}}^{\top}$. If the norm of ${\bm{U}}$ and ${\bm{V}}$ are small, the new update vanishes faster than the original update, and in reverse, if their norm is large, the update blows up, leading to unstable training.
To be specific, we calculate the spectral norm of ${\bm{U}}{\bm{U}}^{\top}$ and ${\bm{V}}^{\top}{\bm{V}}$ and use this to indicate the maximum scale the matrix can stretch a vector. Fig. 11 shows that the largest singular value can vary significantly, being either much greater than or less than 1, depending on the weight shape (input dimension, rank, number of blocks) and magnitude. Interestingly, very structured FFNs with small ranks or many blocks tend to have smaller spectral norms, while others have larger ones. This corresponds to the phenomenon in Fig. 8 and Fig. 9, where smaller FFN are prone to slower convergence and larger ones to loss spikes.
An alternative intuitive perspective of this issue is to realize that learning with structured matrices has additional redundant degrees of freedom brought by the symmetry. For example, to increase the norm of a feature, it can increase either the corresponding weights ${\bm{U}}$ or ${\bm{V}}$. Given that gradient descent makes this decision independently, it will overshoot and make learning less well-behaved.
@@ A.2 On the poorer training dynamics
Appendix B Supplementary experiment
B.1 Design choice
For the BlockDense, we also investigate the opposite order that places the low-rank or dense matrix first, followed by the block-diagonal matrix. However, changing the order does not result in good performance. For example, it increases the perplexity on the RefinedWeb dataset from 29.17 to 29.65 on Transformer-s.
B.2 Ablation of self-guided training
Table 4: Ablation study of self-guided training indicated with for Structured FFN with 32% parameters. Model FLOPs are calculated on one sample with 1024 sequence length.
Architecture Model FFN Model Training Loss PPL Size (M) Size (M) FLOPs (G) Tokens (B) Steps (K) FLOPs Transformer-s 110 56.62 262.9 2.2 4K 1.69e+18 3.2569 25.97 LowRank 73.95 20.94 189.8 2.2 4K 1.22e+18 3.3748 29.22 LowRank 2.2 4K 1.39e+18 3.3329 28.02 BlockDense 73.95 20.94 189.8 2.2 4K 1.22e+18 3.3731 29.17 BlockDense 2.2 4K 1.39e+18 3.3338 28.04 BlockShuffle 73.95 20.94 189.8 2.2 4K 1.22e+18 3.3994 29.95 BlockShuffle 2.2 4K 1.39e+18 3.3583 28.74 Transformer-m 335.08 201.33 788.7 6.7 13K 1.55e+19 2.9062 18.29 LowRank 202.43 68.68 517.0 6.7 13K 1.01e+19 3.0251 20.60 LowRank 6.7 13K 1.21e+19 2.9907 19.90 BlockDense 198.67 64.91 509.3 6.7 13K 1.00e+19 3.0371 20.85 BlockDense 6.7 13K 1.19e+19 3.0008 20.10 BlockShuffle 202.43 68.68 517.0 6.7 13K 1.01e+19 3.0501 21.12 BlockShuffle 6.7 13K 1.21e+19 3.0135 20.36
@@ B.2 Ablation of self-guided training
Method PPL Transformer-s 25.97 Direct decomposition 28.56 Progressive decreasing rank 29.35 Self-guided training 28.02 Self-guided training (slower) 27.90 Transformer-m 18.29 Self-guided training 19.90 Self-guided training (slower) 19.81
B.3 Vision tasks
Method Structured first FFN Dense first FFN #Params (M) Accuracy #Params (M) Accuracy 5-layer MLP (H=768) 4.14 66.99 4.14 66.99 LowRank (R=192) 1.63 64.04 3.26 65.42 BlockShuffle (B=4) 1.63 67.08 3.26 65.67 ViT (H=384) 21.34 92.49 21.34 92.49 LowRank (R=24) 8.29 89.56 9.38 92.09 BlockShuffle (B=16) 8.29 90.42 9.38 92.49
In Table 7, experiments are conducted on 5-layer MLP (MultiLayer Perceptron) and ViT models. The 5-layer MLP consists of a linear layer, batch normalization [49], and ReLU activation function with a hidden dimension of 384. It is trained for 500 epochs with a learning rate of 1.0e-3 and a batch size of 128. For the ViT models, 12 layers with a hidden dimension of 384 are used, and they are trained for 750 epochs with a learning rate of 6.0e-4 and a batch size of 512.
Since the first layer in vision tasks typically prefers from the locality, especially in a 5-layer MLP where the image pixels are directly flattened into the input, we conducted experiments with and without structuring the first layer as LowRank and BlockShuffle. Both setups, particularly when structuring the first linear layer, show that block-diagonal matrices can be advantageous for vision tasks. Specifically, using a block-diagonal matrix to replace the first layer of the 5-layer MLP model often results in superior performance because the block structure groups neighboring pixels together, compensating for the MLP’s lack of locality. However, structuring the first FFN in ViT can lead to significant accuracy degradation, supporting our decision not to apply structured FFNs to the first layer in the main paper.
B.4 Downstream tasks
We also use the Transformer-xl trained on 25.5B tokens for downstream tasks, including PIQA, HellaSwag, Winogrande, and ARC tasks. The evaluation code is based on lm-evaluation-harness [^4] https://github.com/EleutherAI/lm-evaluation-harness with the default prompt. Note that since we train the largest model on a limited number of tokens compared to the 300B tokens used in recent papers [46, 50], the models may not perform optimally on downstream tasks in a zero-shot setting. However, our aim here is to compare the downstream performance of structured FFN and standard FFN.
From Sec. B.4, it can be seen that with structured FFN using 63% of the parameters, we observe an average accuracy degradation of 0.6 to 0.9 on these downstream tasks. With 32% of the parameters, the accuracy performance decreases by about 1.2 to 1.7 percent. While our model is trained on a limited number of tokens and evaluated with default prompts, we conjecture that training on a large enough dataset or finding prompts more suitable for structured FFN might even boost the performance.
Table 8: Performance of Transformer-xl on downstream tasks under the zero-shot setting, where PPL refers to the validation PPL on RefinedWeb dataset.
Architecture Model Size FFN Size PPL PIQA HellaSwag Winogrande ARC (easy) Avg. Transformer-xl 1274M 805M 12.46 70.46 36.62 51.62 46.30 51.25 63% LowRank (R=1024) 985M 516M 12.86 70.51 35.66 50.36 46.21 50.69 BlockDense (B=4, R=1536) 955M 486M 12.97 69.48 36.01 51.46 45.24 50.55 BlockShuffle (B=2) 985M 516M 12.98 68.61 35.54 50.99 46.34 50.37 32% LowRank (R=512) 744M 275M 13.55 68.66 34.73 50.59 44.23 49.55 BlockDense (B=4, R=768) 728M 260M 13.74 69.04 34.37 52.64 44.15 50.05 BlockShuffle (B=4) 744M 275M 13.81 68.61 33.94 50.75 44.23 49.38
B.5 Complete results table

### v2

@@ 4.2 Scaling analysis
Figure 4: Scaling curves of structured matrices with a linear fit for better illustration. The dense model is trained at its optimal trade-off while we train structured FFNs on the same number of tokens and retain 63% or 32% of the original parameters. 1) Structured matrices have steeper scaling curves with much closer results at larger sizes, showing good scaling behavior. 2) With the same training FLOPs, these curves indicate that structured matrices can have fewer parameters and lower validation loss when the x-axis is further extended.
Based on the Chinchilla scaling law, we train four sizes of Transformer models and then use the same amount of tokens to train the structured alternatives. First, we only make the FFN module structured to build the basic understanding, and retain 63% or 32% of the original parameters. In Fig. 4 and Fig. 1, we apply a linear fit to the scaling points for better illustration and show that all three structured matrices have steeper scaling curves compared to the dense models, indicating the significant potential of highly structured large models. More importantly, by fixing the training FLOPs, they have fewer parameters and eventually achieve very close or even slightly lower loss (e.g., LowRank with 63% parameters). Given their steeper scaling curves, we can also expect noticeably lower loss and fewer parameters for structured parameterizations per FLOP when the x-axis is further extended. Detailed numbers are provided in Table 6 in the appendix, with comparisons among the three structured parameterizations.
Next, the attention module is also structured using GQA [2], resulting in wide and structured networks. This further optimizes the use of training FLOPs, addressing the imbalance caused by structuring only the FFN module, which increases the relative impact of the attention module on the overall architecture. We adopt LowRank as an example, as it demonstrates superior performance compared to the other two approaches in our settings, as demonstrated in Table 6 and Fig. 4. To align the training FLOPs, the wide and structured networks are trained on a larger number of tokens. It can be observed in Table 1 that these models achieve lower perplexity while using much fewer parameters. For instance, the parameter count can be reduced from 729M to 464M without compromising perplexity. Additionally, in terms of maximum throughput, ours models achieve an 8% and 17% boost on Transformer-m and Transformer-l, respectively, compared to the fast GQA.
In conclusion, the structured matrices and the wide and structured networks demonstrate great potential in optimizing training FLOP utilization, achieving lower loss with fewer parameters. Additionally, it is important to note that our scaling curves for the structured matrices are not drawn at their optimal training-compute trade-off, while the baseline is.
Scaling model size: better downstream performance
To further illustrate the potential of structured matrices, we consider the overtraining regime and use LowRank as an example. Specifically, we train four sizes of the dense model on 300B tokens, and build the wide and structured networks upon the design of the dense models by applying LowRank to the FFN and reducing the number of attention heads to make the entire network structured. Then, the well-trained models are evaluated on downstream tasks, including PIQA, HellaSwag, Winogrande, and ARC tasks, using lm-evaluation-harness [^4] https://github.com/EleutherAI/lm-evaluation-harness with the default prompt. Fig. 5 presents the results, displaying the scaling trend across the four tasks (see detailed numbers and additional tasks in Sec. A.4). The wide and structured models demonstrate comparable or superior performance, particularly at larger sizes, solidifying their benefits over dense architectures.
Figure 5: Zero-shot performance on downstream tasks in the overtraining regime. The wide and structured networks are built upon dense ones by applying LowRank to the FFN and reducing the number of attention heads to make the entire network structured.
4.3 Efficiency study
Figure 6: Latency of structured and dense FFNs across different FFN widths. Results are evaluated on 30000 tokens. The intermediate size of the FFN is set to be 4 times the FFN width.
Table 2: Training time of Transformer-xl and structured counterparts with 32% and 63% FFN parameters.
Model Params. Training PPL (M) time (h) Transformer-xl 1274 352.2 12.46 63% LowRank 985 302.2 12.86 BlockDense 955 298.7 12.97 BlockShuffle 985 330.6 12.98 32% LowRank 744 260.2 13.55 BlockDense 728 261.2 13.74 BlockShuffle 744 284.9 13.81
Large number of tokens
Using large $T$, the standard linear layers and our efficient structured parametrizations become computation-bound where FLOPs become a latency bottleneck [16]. This setting mainly concerns training, the prefill phase of inference, and extensive offline decoding. In Fig. 6, we evaluate the latency performance of structured and dense FFNs across different FFN widths with 30K tokens. With parameters and FLOPs reduced to 63% or 32%, the lowrank and BlockDense achieve a 1.4$\times$ or a 2.5$\times$ speed-up, respectively. BlockShuffle offers modest improvements, with 1.1$\times$ and 2.0$\times$ speed-ups for the two cases. We also measure the training time of the whole model in Sec. 4.3, and observe that LowRank with 63% FFN parameters reduces the training time by about 15% with 0.4 increased perplexity, and the one with 32% FFN parameters offers 1.35$\times$ whole training speed-up with 1.1 increased perplexity.
Small number of tokens
FFN can be parallelism-bound with small $T$ (e.g., $T=16$) on the A100 GPUs. Then, when $T$ gets increased, FFN becomes memory-bound and will eventually be computation-bound. Online and offline decoding stages may encounter a small number of tokens when unrolling the model step by step. As discussed earlier, our pre-merge method can alleviate the parallelism-bound issue and maintain the same latency with dense matrices. Fig. 7 shows the latency results for three different widths, varying the batch of tokens to determine when to use efficient alternatives or choose pre-merged dense matrices. For example, with a 2048-width FFN, it is difficult to fully utilize resources on GPU with limited tokens. The performance improves significantly when using width 5120 and 6144, such as speed improvements of 2.63$\times$ speed-up of LowRank with 32% FFN parameters on $T=2048$ and 2.81$\times$ acceleration of BlockDense with 32% parameters on $T=1536$.
@@ 4.4 Self-guided training
We apply self-guided training during the first half of training to demonstrate its effectiveness. As shown in Sec. 4.4 and Sec. B.4, our method consistently reduces loss across all efficient parametrizations, improving the perplexity by 1.2 for Transformer-s and 0.8 for Transformer-m. Then, to enable a straightforward comparison under the same training FLOPs, we adjust the training steps for self-guided training and repeat those tokens at the end to ensure they’re fully learned by structured matrices. As can be seen in Sec. B.4 and 14(a), 14(b), Fig. 8, this reduces the perplexity gap for Transformer-xl from 1.0, 1.2, and 1.3 to 0.4, 0.5, and 0.6 for LowRank, BlockDense, and BlockShuffle, respectively, under the same training FLOPs and can still enjoy 32% model FLOPs, which can bring about 2.6$\times$ inference speed-up. Additionally, we compare our method with another advanced baseline that trains structured matrices with more tokens, showing that the self-guided training can achieve comparable or superior results even with the same number of tokens.
Table 3: Performance of the three structured parameterizations when applying self-guided training in the first half of training. This increases 25% FFN training FLOPs. For more comparisons, please refer to Sec. B.3.
Figure 8: Comparisons between dense and structured FFNs with 32% parameters under the same training FLOPs. Structured FFNs are trained either with more tokens or through self-guided training to match training FLOPs. The circle size represents model FLOPs.
5 Conclusion
In this paper, we conducted extensive experiments investigating the use of structured matrices to parameterize FFNs in Transformers, with models scaling up to 1.3B parameters on the RefinedWeb dataset. Our primary aim was not to determine which structured matrix performs best, as this can be task-dependent, but to explore their common properties, including scaling, efficiency, and optimization challenges. We found that all of them exhibit steeper scaling curves compared to dense models. Moreover, our proposed methods, such as self-guided training, can enhance the performance across all structured matrices (e.g., LowRank with the novel training strategy achieves a 1.35× inference speed-up with only a 0.4 increase in perplexity). To conclude, we demonstrate that structured matrices can be strong candidates to replace the dense models in architecture design by scaling studies and also reveal the challenges of applying them.
Limitations: BlockDense and BlockShuffle are more complicated than LowRank. In this work, we only explored a limited range of hyperparameter settings of them. However, since these approaches are new, we believe that further performance improvements may be possible by better tuning their hyperparameters. We primarily focused on language modeling with limited vision experiments included in the appendix. Additionally, we did not explore the optimal scaling laws for structured matrices, which may further enhance performance. We also didn’t investigate models in this paper that are comparable to today’s practical LLMs, such as LLaMA-3. This is not only because of the limited computing resources but also because this study is to start investigating structured parameterizations of linear layers in modern LLM architecture training. We hope our findings and solutions about scaling, efficiency, and optimization will push their usage on the industry side and in future work.
Acknowledgments and Disclosure of Funding
We are grateful to Soham De for the insightful discussions and his valuable feedback on our work. We also sincerely thank the anonymous reviewers for their meaningful reviews and suggestions to make this better. We are also grateful to the RCP and IC cluster system administrators at EPFL for their support and assistance, especially during the project deadline. We also thank nimble.ai for their generous gift to CLAIRE lab, which also helped us to fund some of this research.
Appendix
Appendix A Structured matrices
A.1 Design choice
For BlockDense, we also investigate the reverse order of two projections, placing the low-rank or dense matrix first, followed by the block-diagonal matrix. However, this change surprisingly yields worse performance. For instance, on the RefinedWeb dataset, perplexity increases from 29.17 to 29.65 with Transformer-s and 2.2B training tokens. In the case of BlockShuffle, unlike Dao et al. [13], Zhang et al. [14] does not include a second shuffle operation to restore the original order. Taking this into account, we also experimented with removing the second shuffle operation and found almost no impact on performance. For example, with Transformer-s and -m using 32% FFN parameters, BlockShuffle without the second shuffle achieves perplexities of 29.89 and 21.19, respectively, compared to 29.95 and 21.12 for our adopted version. Nonetheless, we maintain the design from Dao et al. [13] for consistency.
Method PPL -s size 25.97 Direct decomposition 28.56 Progressive decreasing rank 29.35 Self-guided training 28.02 Self-guided training (slower) 27.90 -m size 18.29 Self-guided training 19.90 Self-guided training (slower) 19.81
A.2 Results on Refinedweb dataset
For the scaling points in Fig. 4, we provide detailed results in Table 6 for easier comparison. First, all efficient parameterizations approach the baseline as model size increases. For instance, LowRank with 32% of the parameters has a loss gap of $0.08$ to Transformer-xl, whereas the gap is $0.12$ at the scale of Transformer-s. Moreover, LowRank and BlockDense, with 63% of the parameters in the FFN, increase the loss by only 0.02 to 0.04 while reducing total training time by approximately 15% on Transformer-xl and Transformer-l. Additionally, they accelerate training by 1.35$\times$ with only a 1-point increase in perplexity on Transformer-xl with 32% of the parameters.
Although the main focus of the paper is not to compare different structured matrices but to showcase their general properties—including scaling, efficiency, and optimization—we still provide comparisons in the appendix. From Table 6, by controlling the training FLOPs and model size to be the same, LowRank and BlockDense demonstrate better performance than BlockShuffle in our main experiments, showing a 0.8 lower perplexity on Transformer-s and a 0.4 lower perplexity on Transformer-m. We think that for FFNs in language models, BlockShuffle may not be the optimal choice. However, we further compare these approaches on CIFAR-10 dataset, showing that block-diagonal matrices can serve as a good inductive bias in vision tasks Sec. A.3.
A.3 Results on CIFAR-10 dataset
In Table 7, experiments are conducted on 5-layer MLP (MultiLayer Perceptron) and ViT models. The 5-layer MLP consists of a linear layer, batch normalization [55], and the ReLU activation function with a hidden dimension of 384. It is trained for 500 epochs with a learning rate of 1.0e-3 and a batch size of 128. For the ViT models, 12 layers with a hidden dimension of 384 are used, and they are trained for 750 epochs with a learning rate of 6.0e-4 and a batch size of 512.
Since the first layer in vision tasks typically prefers from the locality, especially in a 5-layer MLP where the image pixels are directly flattened into the input, we conducted experiments with and without structuring the first layer as LowRank and BlockShuffle. Both sets of controls, particularly when structuring the first linear layer, demonstrate that block-diagonal matrices can be beneficial for vision tasks. Specifically, replacing the first layer of the 5-layer MLP model with a block-diagonal matrix even yields better performance, as the block structure effectively groups neighboring pixels, compensating for the MLP’s lack of locality. However, applying structured FFNs to the first layer of ViT can lead to significant accuracy degradation, reinforcing our decision not to use structured FFNs in the first layer in the main experiments.
Method Structured first FFN Dense first FFN Model size (M) Accuracy Model size (M) Accuracy 5-layer MLP (H=768) 4.14 66.99 4.14 66.99 LowRank (R=192) 1.63 64.04 3.26 65.42 BlockShuffle (B=4) 1.63 67.08 3.26 65.67 ViT (H=384) 21.34 92.49 21.34 92.49 LowRank (R=24) 8.29 89.56 9.38 92.09 BlockShuffle (B=16) 8.29 90.42 9.38 92.49
A.4 Results on downstream tasks
Table 8: Performance on downstream tasks under the zero-shot setting. We report the perplexity performance of the validation set of RefinedWeb. For all downstream tasks except LAMBADA, we report accuracy results. For LAMBADA, we present the results in an accuracy/perplexity format. Implementation details are put in Appendix C.
@@ A.4 Results on downstream tasks
Supplementary to Fig. 5, Sec. A.4 presents the detailed results of wide and structured networks compared to dense models on downstream tasks. These models were trained on 300B tokens and implementation details can be found in Appendix C.
Appendix B Self-guided training
B.1 Visualization of training dynamics
We provide additional training dynamics curves, including BlockDense and BlockShuffle, in Fig. 10 and Fig. 11, illustrating that these structures are more challenging to train compared to standard linear layers. Specifically, they exhibit greater sensitivity to learning rates and are more prone to loss spikes. To mitigate this, we apply self-guided training in the most challenging cases, which results in improved training dynamics, such as faster convergence without loss spikes.
Furthermore, we report the loss spikes observed in the large Transformer-xl model trained on the RefinedWeb dataset in Fig. 12.
B.2 Explanation of the poor training dynamics
$\Delta{\bm{W}}={\bm{W}}^{\prime}-{\bm{W}}=-lr\cdot{\bm{g}}{\bm{x}}^{\top}.$ (3)
And the one for ${\bm{U}}{\bm{V}}$:
Thus, it can be seen from Eq. (missing) 4 that the projections ${\bm{U}}{\bm{U}}^{\top}$ and ${\bm{V}}^{\top}{\bm{V}}$ can disrupt the gradient ${\bm{g}}{\bm{x}}^{\top}$. If the norms of ${\bm{U}}$ and ${\bm{V}}$ are small, the new update vanishes faster than the original update, and in reverse, if their norm is large, the update blows up, leading to unstable training.
To be specific, we calculate the spectral norm of ${\bm{U}}{\bm{U}}^{\top}$ and ${\bm{V}}^{\top}{\bm{V}}$ and use this to indicate the maximum scale the matrix can stretch a vector. Fig. 13 shows that the largest singular value can vary significantly, being either much greater than or less than 1, depending on the shape of the weight (input dimension, rank, number of blocks) and magnitude. Interestingly, very structured FFNs with small ranks or many blocks tend to have smaller spectral norms, while others have larger ones. This corresponds to the phenomenon in Fig. 10 and Fig. 11, where smaller FFN are prone to slower convergence and larger ones to loss spikes.
An alternative intuitive perspective of this issue is to realize that learning with structured matrices has additional redundant degrees of freedom brought by symmetry. For example, to increase the norm of a feature, it can increase either the corresponding weights ${\bm{U}}$ or ${\bm{V}}$. Given that gradient descent makes this decision independently, it will overshoot and make learning less well-behaved.
@@ B.2 Explanation of the poor training dynamics
B.3 Design choice
@@
B.4 Results on Refinedweb Dataset
Sec. B.4 provides more comprehensive results of self-guided training, supplementing Sec. 4.4. For Transformer-s and -m, we first present the results of applying self-guided training to the first half of training in the second row of each table block, demonstrating its effectiveness across all three structured matrices. Additionally, by controlling for equal training FLOPs as described in Sec. 4.4, we show that with self-guided training, all three structured FFNs with 32% parameters incur only a 0.4-0.6 increase in perplexity compared to the baseline, while benefiting from a smaller memory footprint and faster speed at inference time. To further illustrate this, we plot these points with the same training FLOPs in Fig. 14, and specifically for LowRank in Fig. 8, highlighting that self-guided training achieves comparable performance to training on more tokens.